A mass of 0.28 kg is fiixed to the end of a 0.93 m long string that is fixed at the other end. Iniitially at rest, the mass is made to rotate around the fixed end with an angular acceleration of 4.43 rad/s. After how many revolutions is the cetripetal force acting on the mass 502 N ?
Answer #1

Let the revolutions required is n to reach centripetal force acting on the mass 502 N.

Let the final velocity of the mass is v.

Let the final angular velocity is \omega .

Given that the mass m=0.28 kg.

The length of the string l=0.93 m

The angular acceleration is \alpha =4.43 rad/s2.

Now we from centripetal force formula

  F=\frac{mv^{2}}{l}

  Fl от ,2 U* m

     or\: \: \: \: \: \: \: v=\sqrt{\frac{Fl}{m}}

or\: \: \: \: \: \: \: v=\sqrt{\frac{502\times 0.93}{0.28}}

\therefore \: \: \: \: \: \: \: v=40.83328474\: \: \: \: \: m/s

Now the final angular speed is

\omega =\frac{v}{l}

or\: \: \: \: \: \: \omega =\frac{40.83328474}{0.93}

  \therefore \: \: \: \: \: \: \omega =43.90675779\: \: \: \: rad/s

Now let the total angular displacement is \Theta .

We know

\omega ^{2}=2\alpha \Theta

or\: \: \: \: \: \: \: \Theta =\frac{\omega ^{2}}{2\alpha }

   or\: \: \: \: \: \: \: \Theta =\frac{\left ( 43.90675779 \right ) ^{2}}{2\times 4.43 }

\therefore \: \: \: \: \: \: \: \Theta =217.5850315\: \: \: \: rad

Now we know

\Theta =2\pi n

or\: \: \: \: \: \: \: n=\frac{\Theta }{2\pi }

or\: \: \: \: \: \: \: n=\frac{217.5850315 }{2\pi }

\therefore \: \: \: \: \: \: \: n=34.62973331\approx 34.63

\therefore After 34.63 revolutions the centripetal force acting on the mass is 502 N.

In case of any doubt please comment...............